OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 0) and (2, 8)?

OpenStudy (anonymous):

y - y1 = m(x-x1)

OpenStudy (anonymous):

m= (8-0)/ (2-(-2)) =2

OpenStudy (anonymous):

I got y-8=2(x-2) but these are the answer choices y = 2(x + 2) y = 2(x – 2) y = –2(x – 2) y = –2(x + 2)

OpenStudy (anonymous):

1st one

OpenStudy (anonymous):

u have got correctly...do the simplification

OpenStudy (anonymous):

y-8=2(x-2) add 8 on bith sides then y = 2x +4 y=2(x+2)